How many gallons of milk containing 5% butterfat must be mixed with 90 gallons of 1% milk to get 2% milk?
Answer by stanbon(75887)

How many gallons of milk containing 5% butterfat must be mixed with 90 gallons of 1% milk to get 2% milk?
-----------------------------
Equation:
fat + fat = fat
0.05x + 0.01*90 = 0.02(x+90)
Multiply thru by 100 to get:
5x + 1*90 = 2x + 2*90
3x = 90
x = 30 gallons (amount of 5% milk needed for the mixture)
===========================
Cheers,
Stan H.